OpenStudy (anonymous):

what are the x and y intercepts of, y=-2x+10 THANK YOU:)))

OpenStudy (anonymous):

show the steps

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

To find the "y" intercept, you must substitute "x" with 0. So... y=-2(0)+10 y=10 then, to find the "x" intercept, you must substitute "y" with 0. So... 0=-2x+10 2x=10 x=5

OpenStudy (anonymous):

the y intercept is 10 and the x intercept is 5

OpenStudy (anonymous):

o ok so you divide 10 by 2 coorect

OpenStudy (anonymous):

Yessir.

OpenStudy (saifoo.khan):

is bailey a "sir" ?

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

no i am a girl

OpenStudy (anonymous):

Oh... whoops... didn't even see the name... o.O

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

My bad...

OpenStudy (anonymous):

haha it is ok

OpenStudy (anonymous):

Well, I'm glad I could help, now back to math homework of my own XD

OpenStudy (anonymous):

sounds good to me